# SRD: Reinforcement-Learned Semantic Perturbation for Backdoor Defense in VLMs

### v1

@@ 4.3 Leveraging Red Perturbations for Attention Recalibration
Building upon the work of Shtedritski shtedritski2023does et al. , which demonstrated that CLIP radford2021learning models exhibit emergent behavior when exposed to red circle visual prompts, we extend this finding by incorporating red block masks as a perturbation mechanism into our defense framework. Our core idea is that since the attention mechanism in multimodal models is usually non-trivial, we are unable to optimize its attention region directly, and therefore introduce discrete perturbation actions (e.g., red block masking) as an indirect intervention to guide the model to “divert attention” from the trigger region. The compatibility score $s(i,t)$ between an image $i$ and a text query $t$ is computed as: $s(i_{\text{red}},t)$. When a red block mask is applied to an image, altering it to $i_{\text{mask}}$, the model’s output changes as:
@@
We utilize this property to design a Semantic Reward Defense (SRD) framework that takes advantage of deep Q-Networks (DQN) to select the most defensive image regions to apply perturbations. The effectiveness of this method can be mathematically captured by the equation:
$\hat{s}(q,a|i_{\text{mask}},t)=\arg\max_{a}s(i_{\text{mask}},t).$ (7)
@@
Here, $q$ denotes the query (e.g., the identifier of an image region), $a$ represents a candidate answer (e.g., a location or object within the image), and $i_{mask}$ refers to the image modified with a red block mask. Since the perturbation actions are discrete and non-differentiable, we employ reinforcement learning (RL) to train the agent model to autonomously learn an intervention policy that disrupts the activation pathway without relying on any prior knowledge of the backdoor, thereby effectively interrupting potential attention entanglement in poisoned samples and enhancing the model’s robustness.
4.4 Optimizing Intervention Policies with Semantic Rewards
To achieve robust defense against stealthy backdoor attacks in image captioning, we propose a reinforcement learning–based perturbation strategy called Semantic Reward Defense (SRD). The approach combines semantic region sensitivity with deep Q-network (DQN) strategy learning, which guides the intelligentsia through semantic rewards to select perturbation regions that interrupt the association between the trigger and the target semantics while maximizing the preservation of the semantic integrity of the image and the naturalness of the textual description.
We formulate the problem as a discrete reinforcement learning environment. At each time step $t$, the environment state $s_{t}$ represents the current perturbation status of the image, and the action $a_{t}$ corresponds to selecting a specific spatial region and perturbation size. The action space is predefined and consists of red-colored masks with sizes from the discrete set $\{20,40,60,80\}$. These perturbations are placed over the image to influence the semantic perception of the CLIP model, which is known to be particularly sensitive to red-highlighted areas. This allows the agent to interfere with the potential trigger effect even without precise localization of the attack region.
where $\mathcal{S}_{0}$ and $\mathcal{F}_{0}$ are the semantic and fluency scores from the unperturbed version. The reward logic is as follows: (1) If both $\mathcal{S}(I_{t})-\mathcal{S}_{0}\geq\alpha$ and $\mathcal{F}(I_{t})-\mathcal{F}_{0}\geq\beta$, then $r_{t}=3$; (2) If only one improves beyond the threshold, then $r_{t}=1$ or $2$; (3) If either metric decreases significantly, a penalty $r_{t}=-1$ or $-2$ is applied. The design essentially encourages policy learning to target perturbation behaviors that disrupt erroneous semantic associations (induced by backdoor triggers) while ensuring that the semantic and linguistic quality of the generated subtitles is not compromised.
After training, the learned strategy is applied to each training sample to generate a red mask for masking potential triggering areas. The images processed in this way constitute an SRD-enhanced dataset, which is then used to train VLMs. In this process, the connection between the trigger pattern and the target semantics is effectively interrupted, thereby suppressing backdoor activation and improving the robustness of the model.
5 Experiments
5.1 Experiments Setup
Victim models. We adopt OpenFlamingo awadalla2023openflamingo as the victim model, using the 3B version with CLIP ViT-L/14 as the visual encoder. While the language model varies by scale, the vision encoder remains consistent. The 3B model features a lightweight autoregressive language model optimized for multimodal tasks.
Datasets. In our experiments, we used two datasets: COCO lin2014microsoft and Flickr30k young2014image . During the fine-tuning phase, we utilized the COCO dataset, while during the inference phase, both COCO and Flickr30k datasets were used. For the COCO dataset, we selected a subset of 5000 images and their corresponding captions for training and evaluation. Each image in COCO is paired with 5 captions describing the content of the image. To adapt the data for the VLM, we used a unified text prompt: “a photo of” as the initial input. During inference, we also used the test split of the Flickr30k dataset, which contains 1,000 images, each associated with 5 captions.
Attack Configurations. To evaluate our defense strategy, we experiment with six representative backdoor attacks, categorized into trigger-based (BadNet gu2019badnets , TrojVLM, VLOOD) and global modification attacks (Blended chen2017targeted , Shadowcast, VL-Trojan). All models are fine-tuned for up to 20 epochs using the AdamW optimizer with cosine annealing and a 10% poisoning rate. Further implementation details are provided in Appendix A.
Defense Configurations. Given the lack of defense methods specifically designed for captioning in vision-language models (VLMs), we evaluate our strategy against three state-of-the-art defenses originally proposed for image classification: ABL li2021anti , VDC zhu2023vdc , and CT qi2023towards . We also include a no-defense baseline for comparison. Detailed implementation settings are provided in Appendix A.
Evaluation Metrics. We evaluate using three metrics: (1) BLEU@4 and CIDEr for text quality (Section  2), (2) Attack Success Rate (ASR) for attack effectiveness, and (3) Semantic Fidelity Score (SFS) for quality degradation (Section  4.1).
5.2 Main Results
Effectiveness of SRD. Table  2 presents the defense effectiveness of our proposed method, SRD, against six different attack methods evaluated on the COCO and Flickr30k datasets. The No Defense section of the table shows the performance of the attack methods, while the SRD section demonstrates the defense effectiveness of our proposed method. From the table 2, it is evident that all six attack methods are effective, with Attack Success Rates (ASR) reaching approximately 99% for each attack. However, the SRD defense method successfully mitigates the impact of these attacks, reducing the ASR to below 60% across all six attacks on the COCO dataset. Notably, for TrojVLM and Shadowcast, SRD reduces the ASR to as low as 3.6% and 5.6%, respectively. These results indicate that SRD effectively disrupts the activation of the backdoor trigger, achieving strong mitigation across diverse attack types.
Table 2: The defense effectiveness of the proposed SRD on the backdoor model, as well as the attack results without defense. B@4 denotes BLEU-4, and C stands for CIDEr.
@@ 5.2 Main Results
Comparison to Existing Defenses. Table  3 compares our proposed defense method, SRD, with several state-of-the-art defenses under six different backdoor attacks. The No Defense section represents the performance of the attacks without any defense mechanisms. For BadNet and Blended, which are not specifically designed for VLMs, SRD significantly outperforms other defense methods. On the COCO dataset, SRD reduces the ASR to 45.7% and 57.4%, respectively, while other defenses fail to mitigate these attacks effectively, with ASR remaining around 90%. For attacks tailored to VLMs, TrojVLM, Shadowcast, and VL-Trojan, some existing defenses show partial effectiveness. However, SRD consistently achieves lower ASR across most of these attacks, maintaining ASR below 35%. Overall, SRD offers robust and consistent defense performance across a diverse set of attack types, outperforming existing methods both in reducing ASR and preserving text quality.
Table 3: The comparison between the defense methods and the proposed SRD defense, as well as the performance of the backdoor model without defense. C stands for CIDEr.
@@
5.3 Ablation Studies
Effectiveness of RL models. We conducted an ablation study to evaluate the impact of different reinforcement learning (RL) models on our defense method under a fixed 10% poisoning rate. Specifically, we compared random masking, PPO, SAC, and DQN, with results shown in Table 4. Random masking already lowers ASR across all attacks, especially in TrojVLM (ASR drops to 3.8%), suggesting that disturbing trigger locations is helpful. However, it performs worse than RL-based methods, leading to higher ASR and lower CIDEr and SFS scores. Among the RL strategies, DQN achieves the best overall balance between defense and caption quality. Interestingly, SAC performs best against VLOOD, reducing ASR to 0.9%, indicating its suitability for certain attack patterns.
Table 4: The impact of different reinforcement learning (RL) models on the proposed SRD defense, along with the backdoor attack performance without defense.
@@ 5.3 Ablation Studies
Blended 1.4 25.9 99.5 32.9 37.5 70.9 39.4 41.5 63.6 41.9 40.0 59.4 42.5 44.6 57.4
COCO TrojVLM 71.8 41.4 97.6 92.7 65.9 3.8 56.2 47.1 42.3 92.9 66.8 3.5 83.0 69.7 3.6
@@
Effectiveness with Different Poisoning Rate. We further investigated how the performance of our defense method SRD is affected by different poisoning rates in Figure  3. Specifically, we evaluated the defense effectiveness under three poisoning rates: 1%, 5%, and 10%. Our findings show that the defense effectiveness varies with the poisoning rate, and the optimal performance is not always observed at the highest poisoning level. For BadNet, SRD achieves the best performance when the poisoning rate is 1%, with ASR remaining the lowest among the three settings. This indicates that SRD is particularly effective in mitigating low-intensity attacks for this type. These results suggest that the defense performance of SRD can vary across poisoning intensities and attack types. In general, lower poisoning rates tend to yield stronger defense outcomes, likely due to the reduced entrenchment of the trigger pattern within the model.
Figure 3: The performance of our SRD defense method under different poisoning rates.
6 Conclusion
@@
We propose Semantic Reward Defense (SRD), a novel defense against backdoor attacks in vision-language models. SRD combines a DQN-based strategy with the Semantic Fidelity Score (SFS) to guide image perturbations that disrupt attacks without needing trigger knowledge. A red masking mechanism further enhances robustness by leveraging CLIP’s sensitivity to red regions. Experiments show SRD effectively lowers attack success while preserving caption quality, highlighting its potential as a semantic-driven, trigger-agnostic defense. Please see limitations in Supplementary.

### v2

@@ Introduction
Building upon the work of Shtedritski (shtedritski2023does), which demonstrated that CLIP (radford2021learning) exhibits emergent behavior when exposed to red circle visual prompts, we extend this finding by incorporating red block masks as a perturbation mechanism into our defense framework. Our core idea is that since the attention mechanism in multimodal models is usually non-trivial, we are unable to optimize its attention region directly, and therefore introduce discrete perturbation actions (red block masking) as an indirect intervention to guide the model to ”divert attention” from the trigger region. The compatibility score $s(i,t)$ between an image $i$ and a text query t reflects the model’s perception of how well the visual content corresponds to the given textual description. This score is sensitive to changes in the image input. When a red block mask is applied to the image, modifying it to $i_{mask}$, the resulting score typically changes as:
@@
We utilize this property to design a SRD framework that takes advantage of DQN to select the most defensive image regions to apply perturbations. The effectiveness of this method can be mathematically captured by the equation:
$a^{*}=\arg\max_{a}s(i_{mask}(a),t)$ (8)
@@
Here, $a$ represents a candidate answer (a location within the image), and $i_{mask}$ refers to the image modified with a red block mask. By selecting the action $a$ that maximizes $s(i_{mask},t)$, the agent effetively learns to apply perturbations in a way that suppresses the influence of potential backdoor triggers while preserving semantic alignment. Given the discrete and non-differentiable nature of perturbation actions, we employ reinforcement learning to train an agent that learns an intervention policy without relying on any prior knowledge of backdoors.
@@
Table 1: The defense effectiveness of the proposed SRD on the backdoor model, as well as the attack results without defense. B@4 denotes BLEU-4, and C stands for CIDEr. All results are shown in %.
Optimizing Intervention Policies with Semantic Rewards
SRD integrates semantic region sensitivity with DQN-based strategy learning. Guided by semantic rewards, the agent selects perturbation regions that block the link between triggers and target semantics. At the same time, it preserves the core image content and keeps the text description natural. We formulate the problem as a discrete reinforcement learning environment.
At each time step $t$, the state represents the current perturbation status of the image. The action $a_{t}$ corresponds to selecting a spatial position and a red-colored mask of a specific size to insert. The action space is predefined and consists of red-colored masks with sizes from the discrete set $\{20,40,60,80\}$. These perturbations are placed over the image to influence the semantic perception of the CLIP model. This allows the agent to interfere with the potential trigger effect even without precise localization of the attack region.
where $\mathcal{S}_{0}$ and $\mathcal{F}_{0}$ are the semantic and fluency scores from the unperturbed version. The reward logic is as follows: (1) If both $\mathcal{S}(I_{t})-\mathcal{S}_{0}\geq\lambda$ and $\mathcal{F}(I_{t})-\mathcal{F}_{0}\geq\beta$, then $r_{t}=3$ where we set $\lambda=0.1$ and $\beta=0.2$ in our experiments; (2) If only one improves beyond the threshold, then $r_{t}=1$ or $2$; (3) If either metric decreases significantly, a penalty $r_{t}=-1$ or 2-2 is applied. The design essentially encourages policy learning to target perturbation behaviors that disrupt erroneous semantic associations (induced by backdoor triggers) while ensuring that the semantic and linguistic quality of the generated subtitles is not compromised.
@@
Table 2: The comparison between the defense methods and the proposed SRD defense, as well as the performance of the backdoor model without defense. C stands for CIDEr. All results are shown in %.
After training, the SRD is applied to each sample in the fine-tuning dataset of the backdoored model to perform data cleansing. Specifically, it generates a red mask to occlude potential trigger regions in the image. The cleaned samples are then used to fine-tune the VLMs.
@@
Blended 1.4 30.6 99.5 32.9 37.5 70.9 39.4 41.5 63.6 41.9 40.0 59.4 42.5 44.6 57.4
COCO TrojVLM 82.9 49.0 99.8 92.7 65.9 3.8 56.2 47.1 42.3 92.9 66.8 3.5 83.0 69.6 3.6
@@
Table 3: The impact of different RL models on the SRD defense, along with the backdoor attack performance without defense. All results are shown in %.
Experiments
Experiments Setup
Victim Models. We adopt OpenFlamingo (awadalla2023openflamingo) as the victim model, using the 3B version with a fixed CLIP ViT-L/14 visual encoder and a scalable autoregressive language model. This architecture is representative among VLMs.
Datasets. In our experiments, we used two datasets: COCO (lin2014microsoft) and Flickr30k (young2014image). We fine-tuned on COCO and conducted inference on both datasets. From COCO, we selected 5000 images with their 5 captions and used the unified prompt “a photo of” for VLM adaptation. For inference, we additionally used the Flickr30k test split, which includes 1,000 images, each with 5 captions.
Attack Configurations. To evaluate our defense strategy, we experiment with six representative backdoor attacks, categorized into trigger-based (BadNet (gu2019badnets), TrojVLM, VLOOD, VL-Trojan) and global modification attacks (Blended (chen2017targeted), Shadowcast). All models are fine-tuned for up to 20 epochs using the AdamW optimizer with cosine annealing and a 10% poisoning rate.
Defense Configurations. Given the lack of defense methods specifically designed for captioning in VLMs, we evaluate our strategy against three SOTA defenses originally proposed for image classification: ABL (li2021anti), VDC (zhu2023vdc), and CT (qi2023towards). We also include a no-defense baseline for comparison.
Evaluation Metrics. We evaluate using three metrics: (1) BLEU@4 and CIDEr for text quality, (2) Attack Success Rate (ASR) for attack effectiveness, and (3) Semantic Fidelity Score (SFS) for quality degradation.
Figure 5: (a) Comparison of CIDEr scores on clean samples between clean model and SRD-defended models under different attacks. (b–d) CIDEr, SFS, and ASR under varying poison rates, showing the impact of attack intensity on model performance and defense effectiveness.
Main Results
Effectiveness of SRD. Table 1 presents the defense effectiveness of SRD against six attack methods on the COCO and Flickr30k datasets. The “No Defense” results show that all attacks are highly effective, with ASR reaching about 99%. From the Table 1, it is evident that all six attack methods are effective, with ASR reaching approximately 99% for each attack. However, the SRD defense method successfully mitigates the impact of these attacks, reducing the ASR to below 60%. Notably, for TrojVLM and Shadowcast, SRD reduces the ASR to as low as 3.6% and 5.6%, respectively. These results indicate that SRD effectively disrupts the activation of the backdoor trigger, achieving strong mitigation across diverse attack types.
Comparison to Existing Defenses. Table 2 compares SRD with several SOTA defenses under five different backdoor attacks. The No Defense represents the performance of the attacks without any defense mechanisms. For BadNet and Blended, which are not specifically designed for VLMs, SRD significantly outperforms other defense methods. On the COCO dataset, SRD reduces the ASR to 45.7% and 57.4%, respectively, while other defenses fail to mitigate these attacks effectively, with ASR remaining around 90%. For attacks tailored to VLMs, TrojVLM, Shadowcast, and VL-Trojan, some existing defenses show partial effectiveness. However, SRD consistently achieves lower ASR across most of these attacks, maintaining ASR below 35%. In general, SRD offers robust and consistent defense performance across a diverse set of attack types, outperforming existing methods to reduce ASR and preserve text quality.
Impact of Defense on Clean Image Performance. To evaluate whether SRD affects the model’s utility on clean samples, we measure CIDEr scores under various attack settings using clean inputs only. As shown in Figure 5 (a) the clean baseline model is trained and evaluated entirely on clean (non-poisoned) data, and achieves a CIDEr score of 103.9% on the clean test set. After applying SRD to remove backdoors from poisoned models, we evaluate their performance on the clean test set. The resulting CIDEr scores range from 88.6% to 101.2%, with an average of 91.8%, depending on the attack. Importantly, the performance drop remains within 15% in all cases, indicating that our defense largely preserves model utility. The largest degradation (14.7%) occurs under VL-Trojan, likely due to stronger entanglement between trigger and task-relevant features. Nevertheless, the model still retains acceptable performance. These results confirm that SRD removes backdoors effectively while maintaining high performance on clean data, making it suitable for real-world use.
Ablation Studies
Effectiveness of RL models. We conducted an ablation study to evaluate the impact of different RL models on our defense method under a fixed 10% poisoning rate. Specifically, we compared random masking, PPO (schulman2017proximal), SAC (haarnoja2018soft), and DQN, with results shown in Table 3. Random masking already lowers ASR across all attacks, especially in TrojVLM (ASR drops to 3.8%), suggesting that disturbing trigger locations is helpful. However, it performs worse than RL-based methods, leading to higher ASR and lower CIDEr and SFS scores. Among the RL strategies, DQN achieves the best overall balance between defense and caption quality. Interestingly, SAC performs best against VLOOD, reducing ASR to 0.9%, indicating its suitability for certain attack patterns.
@@
Effectiveness with Different Poisoning Rate. We further investigated how the performance of SRD is affected by different poisoning rates in Figure 5. Specifically, we evaluated the defense effectiveness under three poisoning rates: 1%, 5%, and 10% in Figure 5 (b-d). Our findings show that the defense effectiveness varies with the poisoning rate, and the optimal performance is not always observed at the highest poisoning level. For BadNet, SRD achieves the best performance when the poisoning rate is 1%, with ASR remaining the lowest among the three settings. This indicates that SRD is particularly effective in mitigating low-intensity attacks for this type. These results suggest that the defense performance of SRD can vary across poisoning intensities and attack types. In general, lower poisoning rates tend to yield stronger defense outcomes, likely due to the reduced entrenchment of the trigger pattern within the model.
Conclusion
@@
We propose Semantic Reward Defense (SRD), a novel defense against backdoor attacks in vision-language models. SRD combines a DQN-based strategy with the Semantic Fidelity Score (SFS) to guide image perturbations that disrupt attacks without needing trigger knowledge. Experiments show SRD effectively lowers attack success while preserving caption quality, highlighting its potential as a semantic-driven, trigger-agnostic defense.
Limitation. Red masking may suppress essential visual cues, and SFS depends on current language models, which may overlook subtle semantics.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant Nos. U23A20318 and 62276195), the Foundation for Innovative Research Groups of Hubei Province (Grant No. 2024AFA017), the Science and Technology Major Project of Hubei Province (Grant No. 2024BAB046), the program “Excellence initiative – research university” for the AGH University of Krakow, as well as the ARTIQ project UMO-2021/01/2/ST6/00004 and ARTIQ/0004/2021, and by funds from the Polish Ministry of Science and Higher Education assigned to the AGH University of Krakow. Dr Tao’s research is partially supported by NTU RSR and Start Up Grants.